When a single card is drawn from a shuffled deck of cards, find the odds against getting a 9.
Answer by Alan3354(21581)

When a single card is drawn from a shuffled deck of cards, find the odds against getting a 9.
--------
There are 4 9's, so the odds of getting a 9 are 4/52 = 1/13.
The odds against are 1 - 1/13 = 12/13